Differentiate the following with respect to x

e^x log x/ x+1
1
Expert's answer
2020-09-24T17:54:08-0400

Not sure whether it is .../(x+1) or .../x + 1 so I'll just provide both.

1.


"(\dfrac{e^x\log x}{x+1})' = \\
=\dfrac{(e^x\log x)'\cdot (x+1)-e^x\log x\cdot (x+1)'}{(x+1)^2}=\\
=\dfrac{(e^x\log x+\frac{e^x}{x})\cdot (x+1)-e^x\log x}{(x+1)^2}=\\
=\dfrac{xe^x\log x+e^x+\frac{e^x}{x}}{(x+1)^2}=e^x\cdot\dfrac{x\log x+1+\frac{1}{x}}{(x+1)^2}"

2.


"(\dfrac{e^x\log x}{x}+1)' = \\
=\dfrac{(e^x\log x)'\cdot x-e^x\log x\cdot (x)'}{x^2}=\\
=\dfrac{(e^x\log x+\frac{e^x}{x})\cdot x-e^x\log x}{x^2}=\\
=\dfrac{xe^x\log x+e^x-e^x\log x}{x^2}=\\
=e^x\cdot\dfrac{x\log x+1-\log x}{x^2}"